# Topological classes of thermodynamics of the rotating charged AdS black holes in gauged supergravities

### v2

@@ IV Five-dimensional rotating charged AdS black holes in minimal gauged supergravity theory
$\phi^{r_{h}}$ $=$ $\frac{\pi\bigg{\{}\Big{\{}3(q+ab)^{2}-r_{h}^{4}\Big{[}3+4\pi P(a^{2}+b^{2})\Big{]}-8\pi Pr_{h}^{6}\Big{\}}\Big{[}\Xi_{a}\Xi_{b}-2(\Xi_{a}+\Xi_{b})\Big{]}\tau-6\pi\Xi_{a}\Xi_{b}r_{h}\Big{[}3r_{h}^{4}+(a^{2}+b^{2})r_{h}^{2}-ab(ab+q)\Big{]}\bigg{\}}}{12r_{h}^{3}\Xi_{a}^{2}\Xi_{b}^{2}\tau}\,,\quad$
@@
$\tau=\frac{6\pi r_{h}(4\pi Pa^{2}-3)(4\pi Pb^{2}-3)\Big{[}ab(ab+q)-r_{h}^{2}(a^{2}+b^{2})-3r_{h}^{4}\Big{]}}{\bigg{\{}4\pi P\Big{[}a^{2}(4\pi Pb^{2}+3)+3b^{2}\Big{]}-27\bigg{\}}\bigg{\{}8\pi Pr_{h}^{6}+\Big{[}4\pi P(a^{2}+b^{2})+3\Big{]}r_{h}^{4}-3(ab+q)^{2}\bigg{\}}}$ (21)
@@ V Conclusions
In this work, we have obtained two particularly exciting consequences: (i) the number of rotation parameters has an important effect on the topological numbers of five-dimensional rotating AdS black holes, and (ii) the number of the electric charge parameters does not contribute to the topological numbers of the four- and five-dimensional rotating AdS black holes. A most related issue is to investigate the topological numbers of black holes in some modified gravity theories, e.g., the scalar-tensor-vector gravity 2406.00579 , Kalb-Ramond gravity 2406.13461 , etc.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12205243, No. 12375053, No. 12205032, and No. 12275037, by the Sichuan Science and Technology Program under Grant No. 2023NSFSC1347, by the Program of Chongqing Human Resources and Social Security Bureau under Grant No. cx2021044, by the Doctoral Research Initiation Project of China West Normal University under Grant No. 21E028, and by the Talent Introduction Program of Chongqing University of Posts and Telecommunications under Grant No. E012A2021209.

### v3

@@ III.2 Rotating pairwise-equal four-charge AdS case
III.3 Discussions
In this subsection, we explore the underlying reasons why multiple electric charge parameters do not alter the topological number of four-dimensional rotating AdS black holes. Now, according to Ref. 2402.00106 , one can construct a new vector
$\Psi^{r_{h}}=\frac{\partial\tau}{\partial r_{h}}\,,\qquad\Psi^{\Theta}=-\cot\Theta\csc\Theta\,.$ (14)
Then, by solving the equation $\Psi^{r_{h}}=0$ and taking the limit $r\to 0$, one can obtain the following critical relationship:
$P_{c}=\frac{3}{8\pi a^{2}}\,.$ (15)
Therefore, the presence or absence of generation and annihilation points is independent of the electric charge parameters. Additionally, according to the results given in our previous work PRD107-084002 , the generation and annihilation points of a four-dimensional Kerr-AdS black hole appear or disappear simultaneously in pairs, thereby not changing its topological number. These conclusions also hold for rotating charged AdS black holes in four-dimensional gauged supergravity theories.
@@ IV Five-dimensional rotating charged AdS black holes in minimal gauged supergravity theory
$\phi^{r_{h}}$ $=$ $\pi(12r_{h}^{3}\Xi_{a}^{2}\Xi_{b}^{2}\tau)^{-1}\bigg{\{}\Big{\{}-r_{h}^{4}\Big{[}3+4\pi P(a^{2}+b^{2})\Big{]}$
$-8\pi Pr_{h}^{6}+3(q+ab)^{2}\Big{\}}\Big{[}\Xi_{a}\Xi_{b}-2(\Xi_{a}+\Xi_{b})\Big{]}\tau$
$-6\pi\Xi_{a}\Xi_{b}r_{h}\Big{[}3r_{h}^{4}+(a^{2}+b^{2})r_{h}^{2}-ab(ab+q)\Big{]}\bigg{\}}\,,\quad$
@@
$\tau$ $=$ $6\pi r_{h}(4\pi Pa^{2}-3)(4\pi Pb^{2}-3)\Big{[}ab(ab+q)-r_{h}^{2}(a^{2}+b^{2})$ (23)
$-3r_{h}^{4}\Big{]}\bigg{\{}4\pi P\Big{[}a^{2}(4\pi Pb^{2}+3)+3b^{2}\Big{]}-27\bigg{\}}^{-1}\bigg{\{}8\pi Pr_{h}^{6}$
$+\Big{[}4\pi P(a^{2}+b^{2})+3\Big{]}r_{h}^{4}-3(ab+q)^{2}\bigg{\}}^{-1}$
@@ IV.3 Double-rotating Kerr-AdS case
IV.4 Discussions
In this subsection, we aim to investigate why multiple electric charge parameters do not affect the topological number of five-dimensional double-rotating AdS black holes. By utilizing the definition of the vector $\Psi$ in Eq. (14) and solving the equation $\Psi^{r_{h}}=0$, two critical pressure values are obtained as
$P_{c_{1}}=\frac{3}{4\pi a^{2}}\,,\qquad P_{c_{2}}=\frac{3}{4\pi b^{2}}\,.$ (24)
Thus, the presence or absence of generation and annihilation points remains unaffected by the electric charge parameters. Moreover, as indicated in Ref. PRD107-084002 , the generation and annihilation points of the five-dimensional singly-rotating Kerr-AdS black hole emerge or vanish in pairs, and thus do not change its topological number. These findings are equally applicable to double-rotating charged AdS black holes in five-dimensional gauged supergravity theories.
@@ V Conclusions
In this work, we have obtained two particularly exciting consequences: (i) the number of rotation parameters has an important effect on the topological numbers of five-dimensional rotating AdS black holes, and (ii) the number of the electric charge parameters does not contribute to the topological numbers of the four- and five-dimensional rotating AdS black holes. A most related issue is to investigate the topological numbers of black holes in some modified gravity theories, e.g., the scalar-tensor-vector gravity 2406.00579 , Kalb-Ramond gravity 2406.13461 ; 2407.07416 , etc.
@@ Acknowledgements.
We are greatly indebted to the anonymous referee for his/her constructive comments to improve the presentation of this work. This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12205243, No. 12375053, No. 12205032, and No. 12275037, by the Sichuan Science and Technology Program under Grant No. 2023NSFSC1347, by the Program of Chongqing Human Resources and Social Security Bureau under Grant No. cx2021044, by the Doctoral Research Initiation Project of China West Normal University under Grant No. 21E028, and by the Talent Introduction Program of Chongqing University of Posts and Telecommunications under Grant No. E012A2021209.